• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name bogdanw

    Log entry time 16:16:16 on April 7,2007

    Entry number 198091

    keyword=shift summary

    investigation of the HRS-L detector: ADC spectra of S1/S2/GC
    and TDC of S1R. The analog signal looks inside the gate.
    TDC reappear after about 30': possibly problem overheating
    when the hut is closed but air-cooling was not ON.
    Ed fixed the AC power.
    Start adjustment of HV on S1 and S2 for HRS-L.
    ALex prepared script for spectra analysis.
    Adjusted HV for S1 and S2 for A_av = 1000-1200
    Start adjustment of HV on GC by using single ph.e
    amplitude as indicator of PMT gain.
    Bob called and told that recently extra delay was
    added to the trigger, so the ADC gate could be late.
    Used 50 ns cable as extra delay for S1l PMT ADC and
    with cosmic events had find that the gate timeing
    is a cause of problem.
    14:45 - restore HV settings
    run 2162 - QE He(e,e'p)
    15:10 - setting for He(e,e'He)
    16:05 - HRS-R dipole is ready